General Physics

on thermometer the freezing point of water is marked as 20° and the boiling point of water is marked as 150°. A temperature of 60°C will be read on this thermometer as
A)110°
B)98°
C)75°
D)40°

Debashish Rout
8 Years agoGrade

1 Answer

Arun
8 Years ago
Let’s determine the difference of the freezing and boiling points. 
d = 150 – 20 = 130 
On this thermometer, there are 130 spaces between the freezing and boiling points of water. On a normal thermometer, the freezing point is 0 and the boiling point is 100. On this thermometer, there are 100 spaces between the freezing and boiling points of water. Let’s use the following equation to determine the fraction of this space that is 60°. 
Fraction = 60/100 = 0.6 
Let’s use the following equation to determine this temperature on the other thermometer. 
Tf = Ti + 0.6 * 130, Ti = 20 
Tf = 20 + 0.6 * 130 = 98° 
Hence option B is correct.
 
Regards
Arun (askIITians forum expert)